Solve:
log(6)(x+7)-log(6)(x-2)=log(6)5

Answer by nerdybill(7384)

Apply "log rules":
log(6)(x+7)-log(6)(x-2)=log(6)5
log(6)((x+7)/(x-2))=log(6)5
(x+7)/(x-2) = 5
(x+7) = 5(x-2)
x+7 = 5x-10
7 = 4x-10
17 = 4x
17/4 = x